OpenStudy (anonymous):

If x^2-4x=4/x-2=2, how can you use the quadratic formula (what does it look like?) to sold for x? tks.

OpenStudy (mathmath333):

\(\large\tt \color{black}{x^2-4x=\dfrac{4}{x-2}=2}\) is it this

OpenStudy (fanduekisses):

No the quadratic formula is \[\frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\]